# Nuclear Toeplitz operators between Fock spaces

### v1

@@ 2. Preliminaries
$\displaystyle\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})^{\prime}$ $\displaystyle\cong\mathcal{N}((F_{\alpha}^{q})^{\prime},(F_{\alpha}^{p})^{\prime})^{\prime}$
$\displaystyle\cong\mathcal{L}((F_{\alpha}^{q})^{\prime\prime},(F_{\alpha}^{p})^{\prime\prime})$
$\displaystyle\cong\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p}).$
The first step is obtained from Theorem 2.3, the second step from Proposition 2.4, the last step follows from the reflexivity.
Let $S=\sum_{j=1}^{\infty}y_{j}\otimes x_{j}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $T\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$, applying Proposition 2.4, we obtain the duality bracket is given by
$\displaystyle\langle S,T\rangle$ $\displaystyle=\langle S^{*},\tau T\tau^{-1}\rangle_{\mathcal{N},\mathcal{L}}$
$\displaystyle=\sum_{j=1}^{\infty}\langle y_{j},\tau Tx_{j}\rangle$
$\displaystyle=\sum_{j=1}^{\infty}\langle Tx_{j},y_{j}\rangle$
$\displaystyle=Tr(TS)$
$\displaystyle=Tr(ST).$
The other proof is similar. ∎
@@
3. Nuclear class Toeplitz operator
In this section, give a complex Borel measure $\mu$, we are going to characterize when the Toeplitz operator $T_{\mu}$ belongs to nuclear class.
Let $1\leq p\leq\infty$ and $p^{\prime}$ to be the conjugate exponent of $p$. For $f\in F_{\alpha}^{p^{\prime}}$, we regard $f$ as an element of $(F_{\alpha}^{p})^{\prime}$ under the pairing $\langle\cdot,\cdot\rangle_{F_{\alpha}^{2}}$, that is
@@ 3. Nuclear class Toeplitz operator
    If $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, and $p<q$, then $\widetilde{T}(z)\in L^{r}(dA)$ for all $(\frac{1}{p^{\prime}}+\frac{1}{q})^{-1}\leq r\leq\infty$.
@@
    If $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, and $q\leq p$, then $\widetilde{T}(z)\in L^{r}(dA)$ for all $1\leq r\leq\infty$.
@@
$\displaystyle\widetilde{T}(z)$ $\displaystyle=\langle Tk_{z},k_{z}\rangle_{F_{\alpha}^{2}}$
$\displaystyle=\sum_{j=1}^{\infty}\langle k_{z},f_{j}\rangle_{F_{\alpha}^{2}}\langle g_{j},k_{z}\rangle_{F_{\alpha}^{2}}$
$\displaystyle=\sum_{j=1}^{\infty}\overline{f_{j}(z)}g_{j}(z)e^{-\alpha|z|^{2}}.$
Let $1\leq r\leq\infty$, we can find $m\geq 1,n\geq s$ such that $\frac{1}{m}+\frac{1}{n}=\frac{1}{r}$. It follows from [19, Theorem 2.10], we have $f_{j}\in F_{\alpha}^{1}\subset F_{\alpha}^{m}$, $g_{j}\in F_{\alpha}^{s}\subset F_{\alpha}^{n}$, and
@@
Then using Hölder’s inequality, we have
$\displaystyle\|\widetilde{T}\|_{L^{r}}$ $\displaystyle\leq\sum_{j=1}^{\infty}\|{f_{j}(z)}e^{-\frac{\alpha}{2}|z|^{2}}\|_{L^{m}}\|g_{j}(z)e^{-\frac{\alpha}{2}|z|^{2}}\|_{L^{n}}$
$\displaystyle\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,m}\|g_{j}\|_{\alpha,n}$
$\displaystyle\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,s}$
$\displaystyle\lesssim\sum_{j=1}^{\infty}\|{f_{j}}\|_{(f_{\alpha}^{\infty})^{\prime}}\|g_{j}\|_{\alpha,s}.$
Now, considering the infimum over all possible representations (2.1) yields
@@
Since $|z-z_{0}|<1$, there exists a constant $C>0$ such that $|(k_{z}(w)-k_{z_{0}}(w))e^{-\frac{\alpha}{2}|w|^{2}}|\leq Ce^{-\frac{\alpha}{4}|w|^{2}}$. According to the dominated convergence theorem, the conclusion holds for $1\leq p<\infty$. When $p=\infty$, the conclusion follows from the fact $\|f\|_{\alpha,\infty}\leq\|f\|_{\alpha,p}$ for any $1\leq p<\infty$. ∎
Lemma 3.4.
Let X be a Banach space, and let $\sum_{n=1}^{\infty}x_{n}$ be a absolutely convergent series in X with sum A, Then any rearrangement $\sum_{n=1}^{\infty}x_{\pi(n)}$ induced by a permutation $\pi:\mathbb{N}\rightarrow{\mathbb{N}}$ also converges to A.
Fix $\epsilon>0$. Choose $N\in\mathbb{N}$ such that:
$\sum_{k=N+1}^{\infty}\|x_{k}\|<\frac{\epsilon}{2}.$
Let $S_{n}=\sum_{k=1}^{n}x_{k}$. There exists $N_{1}\geq N$ such that for all $m\geq N_{1}$:
$\left\|\sum_{n=1}^{m}x_{n}-A\right\|<\frac{\epsilon}{2}.$
Choose $M\in\mathbb{N}$ such that $\{1,\ldots,N_{1}\}\subseteq\{\pi(1),\ldots,\pi(M)\}$. For any $m>M$, Split the rearranged partial sum:
$S^{\prime}_{m}=\sum_{\begin{subarray}{c}k\leq m\\ \pi(k)\leq N_{1}\end{subarray}}x_{\pi(k)}+\sum_{\begin{subarray}{c}k\leq m\\ \pi(k)>N_{1}\end{subarray}}x_{\pi(k)}.$
By the triangle inequality:
$\|S^{\prime}_{m}-A\|\leq\|S_{N_{1}}-A\|+\sum_{k=N_{1}+1}^{\infty}\|x_{k}\|<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon.$
since $\epsilon$ is arbitrary. The rearranged series converges to $A$. ∎
For a fixed positive number $r$, let
$r\mathbb{Z}^{2}=\{nr+\mathrm{i}mr:n,m\in\mathbb{Z}\}$
Then, we define
$S_{r}=\{z=x+\mathrm{i}y:-r/2\leq x<r/2,\;-r/2\leq y<r/2\}.$
Clearly, we have the disjoint decomposition for $\mathbb{C}$:
$\mathbb{C}=\bigcup_{w\in r\mathbb{Z}^{2}}(S_{r}+w),$
Consequently, for every $f\in L^{1}(\mathbb{C},\mathrm{d}\mu)$ we have
$\int_{\mathbb{C}}f(z)\,\mathrm{d}\mu(z)=\sum_{w\in r\mathbb{Z}^{2}}\;\int_{S_{r}+w}f(z)\,\mathrm{d}\mu(z).$
@@
Let $\mu$ be a complex Borel measure on $\mathbb{C}$, and $|\mu|(\mathbb{C})<\infty$. Then $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$. Moreover, we have
@@
Let $z\in\mathbb{C}$, we consider the operator $k_{z}\otimes k_{z}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$, i.e.
@@
On the one hand, there exists an interger $N_{1}$ such that $r_{N}<\sqrt{2}\delta$ for all $N>N_{1}$. Fix a $N>max(N_{0},N_{1})$, then by Lemma 3.4, for any $N^{\prime}>N$, we have
$\displaystyle T_{r_{N}}-T_{r_{N^{\prime}}}=\sum_{k=1}^{\infty}\mu(S_{r_{N}}+a_{k}^{(N)})k_{a_{k}^{(N)}}\otimes k_{a_{k}^{(N)}}-\sum_{j=1}^{\infty}\mu(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})k_{a_{j}^{(N^{\prime})}}\otimes k_{a_{j}^{(N^{\prime})}}$
$\displaystyle=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})](k_{a_{k}^{(N)}}\otimes k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}}\otimes k_{a_{j}^{(N^{\prime})}})$
$\displaystyle=I_{1}+I_{2}$
@@
$\displaystyle I_{1}=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})](k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\otimes k_{a_{j}^{(N^{\prime})}}$
$\displaystyle I_{2}=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}}).$
@@
$\displaystyle\|\sum_{k\in S_{1}}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\mathcal{N}}$
$\displaystyle\leq\sum_{k\in S_{1}}\sum_{j=1}^{\infty}|\mu|[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]\|(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\alpha,1}\|k_{a_{j}^{(N^{\prime})}}\|_{(f_{\alpha}^{\infty})^{\prime}}$
$\displaystyle\leq|\mu|(\mathbb{C})\epsilon.$
@@
$\displaystyle\|\sum_{k\in S_{2}}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\mathcal{N}}$
$\displaystyle\leq 2|\mu|(\mathbb{C}\setminus\overline{B(0,R)})$
$\displaystyle\leq 2\epsilon.$
@@
$\displaystyle|\widetilde{T_{r}}(z)-\frac{\pi}{\alpha}\widetilde{\mu}(z)|=|\sum_{k=1}^{\infty}\mu(S_{r}+a_{k})e^{-\alpha|z-a_{k}|^{2}}-\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)|$
$\displaystyle=|\sum_{k=1}^{\infty}\int_{S_{r}+a_{k}}e^{-\alpha|z-w|^{2}}-e^{-\alpha|z-a_{k}|^{2}}d\mu(w)|$
$\displaystyle\leq\sum_{k=1}^{\infty}\int_{S_{r}+a_{k}}|e^{-\alpha|z-w|^{2}}-e^{-\alpha|z-a_{k}|^{2}}|d|\mu|(w).$
@@
$\displaystyle|\widetilde{T_{r}}(z)-\frac{\pi}{\alpha}\widetilde{\mu}(z)|<\epsilon.$
This implies $lim_{r\rightarrow 0}\widetilde{T_{r}}(z)=\frac{\pi}{\alpha}\widetilde{\mu}(z)$, hence $\widetilde{T}(z)=\frac{\pi}{\alpha}\widetilde{\mu}(z)$.
Set $F(z,w)=T_{\mu}K_{z}(w)$ and $G(z,w)=TK_{z}(w)$. These functions are analytic with respect to $w$, and are anti-holomorphic with respect to $z$. Furthermore, we have $F(z,z)=\frac{\alpha}{\pi}G(z,z)=e^{\alpha|z|^{2}}\widetilde{\mu}(z)$. A standard polarization theorem in several complex variables (see [14]) yields $F(z,w)=\frac{\alpha}{\pi}G(z,w)$ for all $z,w\in\mathbb{C}$. Then we have $T_{\mu}K_{z}=\frac{\alpha}{\pi}TK_{z}$ for all $z\in\mathbb{C}$. By the density of the finite linear span of kernel functions in $f_{\alpha}^{\infty}$, we have $T_{\mu}=\frac{\alpha}{\pi}T\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$.
@@
Recall that $F_{\alpha}^{1}\subset F_{\alpha}^{p}$ for any $p\geq 1$, and the inclusion is continuous. Therefore, we obtain the following corollary.
Corollary 3.6.
Let $\mu$ be a complex Borel measure on $\mathbb{C}$, and $|\mu|(\mathbb{C})<\infty$. Then $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, where $1\leq p<\infty,1\leq q\leq\infty$. Moreover
$\|T_{\mu}\|_{\mathcal{N}}\lesssim|\mu|(\mathbb{C}).$
By the hypothesis and Theorem 3.5, we have $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$. Then we can find $(f_{j})$, $(g_{j})\subset F_{\alpha}^{1}$ such that $T_{\mu}=\sum_{j=1}^{\infty}f_{j}\otimes g_{j}$, and $\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,1}<\infty$. Since $p^{\prime},q>1$, the embedding from $F_{\alpha}^{1}$ to $F_{\alpha}^{p^{\prime}}$ and $F_{\alpha}^{q}$ is continuous. Hence
$\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,p^{\prime}}\|g_{j}\|_{\alpha,q}\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,1}<\infty.$
Then we have $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$, and $\|T_{\mu}\|_{\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})}\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,1}$. Take the infimum over all possible representations of $T_{\mu}$ on the right-hand side of the inequality, we obtain $\|T_{\mu}\|_{\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})}\lesssim\|T_{\mu}\|_{\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})}\lesssim|\mu|(\mathbb{C})$. ∎
In 2020, Fulsche proved that $T_{f}\in\mathcal{N}(F_{\alpha}^{p})$ if $f\in L^{1}(\mathbb{C})$ (see [9, Lemma 2.11]), thus the above corollary can be viewed as a generalization of this result. Moreover, if $\mu$ is a positive measure, the below theorem shows that the condition $\mu(\mathbb{C})<\infty$ is also sufficient for $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ when $q\leq p$.

### v2

@@ 2.2 Nuclear Operator
$\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})^{\prime}$ $\cong\mathcal{N}((F_{\alpha}^{q})^{\prime},(F_{\alpha}^{p})^{\prime})^{\prime}$
$\cong\mathcal{L}((F_{\alpha}^{q})^{\prime\prime},(F_{\alpha}^{p})^{\prime\prime})$
$\cong\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p}).$
The first step is obtained from Theorem 2.3; the second step from Proposition 2.4; and the last step follows from reflexivity.
Let $S=\sum_{j=1}^{\infty}y_{j}\otimes x_{j}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $T\in\mathcal{L}(F_{\alpha}^{q},F_{\alpha}^{p})$; applying Proposition 2.4, we obtain that the duality bracket is given by
$\langle S,T\rangle$ $=\langle S^{*},\tau T\tau^{-1}\rangle_{\mathcal{N},\mathcal{L}}$
$=\sum_{j=1}^{\infty}\langle y_{j},\tau Tx_{j}\rangle$
$=\sum_{j=1}^{\infty}\langle Tx_{j},y_{j}\rangle$
$=\mathrm{Tr}(TS)$
$=\mathrm{Tr}(ST).$
The other proof is similar.
@@
3 Nuclear Toeplitz operator
In this section, given a complex Borel measure $\mu$, we are going to characterize when the Toeplitz operator $T_{\mu}$ belongs to the nuclear class.
Let $1\leq p\leq\infty$ and $p^{\prime}$ be the conjugate exponent of $p$. For $f\in F_{\alpha}^{p^{\prime}}$, we regard $f$ as an element of $(F_{\alpha}^{p})^{\prime}$ under the pairing $\langle\cdot,\cdot\rangle_{F_{\alpha}^{2}}$, that is
@@ 3 Nuclear Toeplitz operator
    If $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $p\leq q$, then $\widetilde{T}(z)\in L^{r}(dA)$ for all $(\frac{1}{p^{\prime}}+\frac{1}{q})^{-1}\leq r\leq\infty$.
@@
    If $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $p>q$, then $\widetilde{T}(z)\in L^{r}(dA)$ for all $1\leq r\leq\infty$.
@@
$\widetilde{T}(z)$ $=\langle Tk_{z},k_{z}\rangle_{F_{\alpha}^{2}}$
$=\sum_{j=1}^{\infty}\langle k_{z},f_{j}\rangle_{F_{\alpha}^{2}}\langle g_{j},k_{z}\rangle_{F_{\alpha}^{2}}$
$=\sum_{j=1}^{\infty}\overline{f_{j}(z)}g_{j}(z)e^{-\alpha|z|^{2}}.$
Let $1\leq r\leq\infty$, we can find $m\geq 1,n\geq s$ such that $\frac{1}{m}+\frac{1}{n}=\frac{1}{r}$. It follows from [MR2934601, Theorem 2.10], we have $f_{j}\in F_{\alpha}^{1}\subset F_{\alpha}^{m}$, $g_{j}\in F_{\alpha}^{s}\subset F_{\alpha}^{n}$, and
@@
Then by Hölder’s inequality, we have
$\|\widetilde{T}\|_{L^{r}}$ $\leq\sum_{j=1}^{\infty}\|{f_{j}(z)}e^{-\frac{\alpha}{2}|z|^{2}}\|_{L^{m}}\|g_{j}(z)e^{-\frac{\alpha}{2}|z|^{2}}\|_{L^{n}}$
$\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,m}\|g_{j}\|_{\alpha,n}$
$\lesssim\sum_{j=1}^{\infty}\|f_{j}\|_{\alpha,1}\|g_{j}\|_{\alpha,s}$
$\lesssim\sum_{j=1}^{\infty}\|{f_{j}}\|_{(f_{\alpha}^{\infty})^{\prime}}\|g_{j}\|_{\alpha,s}.$
Now, taking the infimum over all possible representations (2.4) yields
@@
Since $|z-z_{0}|<1$, there exists a constant $C>0$ such that $|(k_{z}(w)-k_{z_{0}}(w))e^{-\frac{\alpha}{2}|w|^{2}}|\leq Ce^{-\frac{\alpha}{4}|w|^{2}}$. According to the dominated convergence theorem, the conclusion holds for $1\leq p<\infty$. When $p=\infty$, the conclusion follows from the fact $\|f\|_{\alpha,\infty}\leq\|f\|_{\alpha,p}$ for any $1\leq p<\infty$.
For a fixed positive number $r$, let rZ^2={nr+imr:n,m∈Z}. Then, we define S_r={z=x+iy:-r/2≤x<r/2,-r/2≤y<r/2}. Clearly, we have the disjoint decomposition for $\mathbb{C}$: C=⋃_w∈rZ^2(S_r+w),
@@
Let $\mu$ be a complex Borel measure on $\mathbb{C}$ with $|\mu|(\mathbb{C})<\infty$. Then $T_{\mu}$ belongs to both $\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{s})$ and $\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ for $1\leq p<\infty$ and $1\leq q,s\leq\infty$. Moreover,
@@
Proof 6.
We first proof $T_{\mu}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$. Let $z\in\mathbb{C}$, we consider the operator $k_{z}\otimes k_{z}\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$, i.e.
@@
On the one hand, there exists an integer $N_{1}$ such that $r_{N}<\sqrt{2}\delta$ for all $N>N_{1}$. Fix an $N>\max(N_{0},N_{1})$. For any $N^{\prime}>N$, recall that in a Banach space, any absolutely summable sequence is unconditionally summable [MR1342297, Proposition 1.1]. We have,
$T_{r_{N}}-T_{r_{N^{\prime}}}=\sum_{k=1}^{\infty}\mu(S_{r_{N}}+a_{k}^{(N)})k_{a_{k}^{(N)}}\otimes k_{a_{k}^{(N)}}-\sum_{j=1}^{\infty}\mu(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})k_{a_{j}^{(N^{\prime})}}\otimes k_{a_{j}^{(N^{\prime})}}$
$=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})](k_{a_{k}^{(N)}}\otimes k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}}\otimes k_{a_{j}^{(N^{\prime})}})$
$=I_{1}+I_{2}$
@@
$I_{1}=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})](k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\otimes k_{a_{j}^{(N^{\prime})}}$
$I_{2}=\sum_{k=1}^{\infty}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}}).$
@@
$\|\sum_{k\in S_{1}}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\mathcal{N}}$
$\leq\sum_{k\in S_{1}}\sum_{j=1}^{\infty}|\mu|[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]\|(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\alpha,1}\|k_{a_{j}^{(N^{\prime})}}\|_{(f_{\alpha}^{\infty})^{\prime}}$
$\leq|\mu|(\mathbb{C})\epsilon.$
@@
$\|\sum_{k\in S_{2}}\sum_{j=1}^{\infty}\mu[(S_{r_{N}}+a_{k}^{(N)})\cap(S_{r_{N^{\prime}}}+a_{j}^{(N^{\prime})})]k_{a_{k}^{(N)}}\otimes(k_{a_{k}^{(N)}}-k_{a_{j}^{(N^{\prime})}})\|_{\mathcal{N}}$
$\leq 2|\mu|(\mathbb{C}\setminus\overline{B(0,R)})$
$\leq 2\epsilon.$
@@
$|\widetilde{T_{r}}(z)-\frac{\pi}{\alpha}\widetilde{\mu}(z)|=|\sum_{k=1}^{\infty}\mu(S_{r}+a_{k})e^{-\alpha|z-a_{k}|^{2}}-\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)|$
$=|\sum_{k=1}^{\infty}\int_{S_{r}+a_{k}}e^{-\alpha|z-w|^{2}}-e^{-\alpha|z-a_{k}|^{2}}d\mu(w)|$
$\leq\sum_{k=1}^{\infty}\int_{S_{r}+a_{k}}|e^{-\alpha|z-w|^{2}}-e^{-\alpha|z-a_{k}|^{2}}|d|\mu|(w).$
@@
$|\widetilde{T_{r}}(z)-\frac{\pi}{\alpha}\widetilde{\mu}(z)|<\epsilon.$
This implies $\lim\limits_{r\rightarrow 0}\widetilde{T_{r}}(z)=\frac{\pi}{\alpha}\widetilde{\mu}(z)$, hence $\widetilde{T}(z)=\frac{\pi}{\alpha}\widetilde{\mu}(z)$.
Set $F(z,w)=T_{\mu}K_{z}(w)$ and $G(z,w)=TK_{z}(w)$. These functions are analytic with respect to $w$, and are anti-holomorphic with respect to $z$. Furthermore, we have
$F(z,z)=\frac{\alpha}{\pi}G(z,z)=e^{\alpha|z|^{2}}\widetilde{\mu}(z).$
A standard polarization theorem in several complex variables (see [MR1846625]) yields $F(z,w)=\frac{\alpha}{\pi}G(z,w)$ for all $z,w\in\mathbb{C}$. Then we have $T_{\mu}K_{z}=\frac{\alpha}{\pi}TK_{z}$ for all $z\in\mathbb{C}$. By the density of the finite linear span of kernel functions in $f_{\alpha}^{\infty}$, we have $T_{\mu}=\frac{\alpha}{\pi}T\in\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{1})$.
@@
Recall that the normalized reproducing kernel $k_{z}$ is a unit vector in $F_{\alpha}^{p}$ for all $1\leq p\leq\infty$, hence the operator $T_{r}$ defined above belongs to both $\mathcal{N}(f_{\alpha}^{\infty},F_{\alpha}^{s})$ and $\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ for the indices under consideration. Repeating this argument completes the proof for all cases.
From the proof of Theorem 3.4, let
$S=\{T_{\mu}:|\mu|(\mathbb{C})<\infty\},$
viewed as a subset of $\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$. Then we have the inclusion
$S\subset\overline{\operatorname{span}}^{\|\cdot\|_{\mathcal{N}}}\{k_{z}\otimes k_{z}:z\in\mathbb{C}\}.$
In 2020, Fulsche proved that $T_{f}\in\mathcal{N}(F_{\alpha}^{p})$ if $f\in L^{1}(\mathbb{C})$ (see [MR4107813, lemma 2.11], thus the above Theorem can be viewed as a generalization of this result. Moreover, if $\mu$ is a positive measure, Theorem 1.1 shows that the condition $\mu(\mathbb{C})<\infty$ is also sufficient for $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ when $q\leq p$. Next, we give the proof of Theorem 1.1:
Proof 7 (Proof of Theorem 1.1).
A direct calculation shows that $\widetilde{T_{\mu}}(z)=\widetilde{\mu}(z)$. Since $\mu\geq 0$, we can use Fubini’s Theorem to obtain
$\|\widetilde{\mu}\|_{L^{1}}$ $=\frac{\alpha}{\pi}\int_{\mathbb{C}}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}d\mu(w)dA(z)$
$=\frac{\alpha}{\pi}\int_{\mathbb{C}}\int_{\mathbb{C}}e^{-\alpha|z-w|^{2}}dA(z)d\mu(w)$
$=\mu(\mathbb{C}).$
$(1)\Rightarrow(3)$: If $1\leq q\leq p<\infty$, by Theorem 3.2 (3), we have $\widetilde{T_{\mu}}(z)=\widetilde{\mu}(z)\in L^{1}(\mathbb{C},dA)$, which implies $\mu(\mathbb{C})<\infty$. $(3)\Rightarrow(1)$ follows from Theorem 3.4. The proof of $(2)\Leftrightarrow(3)$ is similar.
Estimate (1.1) follows from the equation $\|\widetilde{\mu}\|_{L^{1}}=\mu(\mathbb{C})$, Theorem 3.2 and Theorem 3.4.
Remark 3.5.
Suppose $1\leq p<q,$ and $\mu\geq 0$. If $\mu(\mathbb{C})<\infty$, by Theorem 3.4, we have $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$. However, we cannot use Theorem 3.2 to prove the equivalence between $T_{\mu}\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ and $\mu(\mathbb{C})<\infty$, since in this case $(\frac{1}{p^{\prime}}+\frac{1}{q})^{-1}>1$. In fact, we will show that there exists $T\in\mathcal{N}(F_{\alpha}^{p},F_{\alpha}^{q})$ such that $\widetilde{T}(z)\notin L^{1}$.